OpenStudy (anonymous):

Is the line through points P(3, –5) and Q(1, 4) parallel to the line through points R(–1, 1) and S(3, –3)? Explain. Choose one answer. a. Yes; the lines have equal slopes. b. Yes; the lines are both vertical. c. No; the lines have unequal slopes. d. No; one line has zero slope, the other has no slope.

OpenStudy (anonymous):

you gotta compute the slope of each for example, the slope of the first one is \[\frac{4-(-5)}{1-3}=\frac{9}{-2}\]

OpenStudy (anonymous):

if the second one has the same slope then they are parallel it does not however

OpenStudy (anonymous):

you know how to find the slope of the second one?

OpenStudy (anonymous):

Not really. I've been trying to figure it out. That's why I haven't replied yet.

OpenStudy (anonymous):

it is always \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

for \((-1,1)\) and \((3,-3)\) it would be \[\frac{-3-1}{3-(-1)}\]

OpenStudy (anonymous):

which, if my arithmetic is correct, is just \(-1\)

OpenStudy (anonymous):

Alright I'll try and figure it out. And post it and you tell me if it's right!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is the answer D?

OpenStudy (anonymous):

no D says one line has zero slope, but we calculated the slope of the first one and got \[-\frac{9}{2}\]

OpenStudy (anonymous):

i wrote above how to calculate the slope of the other line can you see where i wrote that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I'm trying to figure out which one i'm calculating.

OpenStudy (anonymous):

you have to calculate both of them to answer this question

OpenStudy (anonymous):

i calculated the first one as \(-\frac{9}{2}\)

OpenStudy (anonymous):

So for the second would I do. 1 - 3 and 4 - (-5) ?

OpenStudy (anonymous):

that is the first one and no that is not quite right you want to do the first or the second you pick we will do it slowly

OpenStudy (anonymous):

Either one is fine.

OpenStudy (anonymous):

ok lets compute the slope of the line with points \((3,-5)\) and \((1,4)\)

OpenStudy (anonymous):

it is always \[\frac{y_2-y_1}{x_2-x_1}\] in your example \(y_2=4,y_1=-5, x_2=1,x_1=3\)

OpenStudy (anonymous):

when we substitute those numbers in to find the slope, you get \[\frac{4-(-5)}{1-3}\]

OpenStudy (anonymous):

after doing the arithmetic, you have a slope of \[-\frac{9}{2}\]

OpenStudy (anonymous):

Okay, I got that finally.

OpenStudy (anonymous):

whew now on to the next one?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

points are \((-1,1)\) and \((3,-3)\)

OpenStudy (anonymous):

it is still \[\frac{y_2-y_1}{x_2-x_1}\] but this time \[y_2=-3,y_1=1,x_2=3,x_1=-1\] you want to try it?

OpenStudy (anonymous):

Yes, please!

OpenStudy (anonymous):

go ahead, let me know what you get

OpenStudy (anonymous):

-1 is what I got.

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

Yay! Finally.

OpenStudy (anonymous):

that means the final answer is C, it is C 95% of the time, so we probably could have guessed that from the beginning

OpenStudy (anonymous):

I was going to say C earlier, but I wasn't sure. can you help me with three more please!

OpenStudy (anonymous):

sure go ahead we can do them

OpenStudy (anonymous):

Thank you! What is an equation in point-slope form for the line perpendicular to y = 2x + 13 that contains (8, –4)? A. y + 4 =-1/2(x-8) B. x + 4 = 2(y – 8) C.y + 8 =-1/2(x-4) D.y + 4 = 2(x – 8)

OpenStudy (anonymous):

i got a raft of question marks instead of copy and paste, it would be best to make a screen shot

OpenStudy (anonymous):

i could try to refresh my browser maybe that would work

OpenStudy (anonymous):

Try and let me know.

OpenStudy (anonymous):

that's better

OpenStudy (anonymous):

It's been doing it to me, as well. So it might be this website.

OpenStudy (anonymous):

perpendicular to \(y=2x+3\) slope of \(y=2x+3\) perpedicular line will have slope \(-\frac{1}{2}\) the "negative reciprocal

OpenStudy (anonymous):

you mean y=2x + 13?

OpenStudy (anonymous):

yeah right makes no difference really

OpenStudy (anonymous):

slope is still \(-\frac{1}{2}\) for the perpendicular line

OpenStudy (anonymous):

Okay gotcha!

OpenStudy (anonymous):

the point is \((8,-4)\) and the slope is \(-\frac{1}{2}\) the point slope formula gives \[y-(-4)=-\frac{1}{2}(x-8)\] or \[y+4=-\frac{1}{2}(x-8)\]

OpenStudy (anonymous):

So would it be A?

OpenStudy (anonymous):

yes, a rare A

OpenStudy (anonymous):

Next Question A. y = 1/3(x) - (-4) B. y= 1/3(x) + (-2/3) C. y= 3(x) + (-10/3) D. 3(x) + (-2/3)

OpenStudy (anonymous):

D. y=*

OpenStudy (anonymous):

this kind of sucks because i cannot read the two points clearly one thing is for sure, the slope is not 3, it is \(\frac{1}{3}\)

OpenStudy (anonymous):

Do you need me to tell you the two points?

OpenStudy (anonymous):

in any case we don't really need to do much more the slope is \(\frac{1}{3}\) because the line goes right 3 up 1 the y intercept is not \(-4\) is is some negative fraction

OpenStudy (anonymous):

from your eyeballs using no computation, you get to pick B

OpenStudy (anonymous):

do you see what i mean by the y intercept? where the line crosses the y axis

OpenStudy (anonymous):

The points are -4, 2 and 5, 1 and yes i see where your coming from.

OpenStudy (anonymous):

-4, -2*

OpenStudy (anonymous):

so we need go no further the y intercept is not -4 pick B

OpenStudy (anonymous):

Okay! Next Question. Marks: 1 What is the equation in point-slope form for the line parallel to y = –2x + 10 that contains J(6, 8)? Choose one answer. a. y – 8 = –2(x – 6) b. y – 8 = 2(x – 6) c. y + 8 = –2(x – 6) d. x – 8 = 2(y – 6)

OpenStudy (anonymous):

You might have to Refresh again. It did the question marks.

OpenStudy (anonymous):

hold on let me refresh again

OpenStudy (anonymous):

ok that's better parallel that means they have the same slope what is the slope of \(y=-2x+10\) ?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

yeah so it is like the other one \[y-y_1=m(x-x_1)\] with \[m=-2, x_1=6,y_1=8\]

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

ok looks good

OpenStudy (anonymous):

Thank you!! One more thing before I submit my test can you check my answers for the questions I did myself?

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

What is an equation for the line that passes through points (4, –4) and (8, 4)? Choose one answer. a. (y + 4) = –2(x – 4) b. (y – 4) = –2(x + 4) c. (y – 4) = 2(x + 4) d. (y + 4) = 2(x – 4) I choose C.

OpenStudy (anonymous):

crap hold on

OpenStudy (anonymous):

ok back

OpenStudy (anonymous):

did you get this \[y+4 = 2 (x-4)\]

OpenStudy (anonymous):

oh no you got the signs backwards

OpenStudy (anonymous):

should be D i think

OpenStudy (anonymous):

What must be true about the slopes of two perpendicular lines, neither of which is vertical? Choose one answer. a. The slopes have product 1. b. The slopes are equal. c. The slopes have product –1. d. One of the slopes must be 0. I choose C again. Write an equation in point-slope form of the line through point J(4, –4) with slope 4. A. y + 4 = 4 (x + 4) B. y - 4 (x + 4) C. y + -4 (x -4) D. y + 4 (x -4) C again.

OpenStudy (anonymous):

Okay I changed it! thank you!

OpenStudy (anonymous):

first one the slopes have a product equal -1 was that the choice C?

OpenStudy (anonymous):

I choose C for both.

OpenStudy (anonymous):

And yes.

OpenStudy (anonymous):

\[y+4=4(x-4)\] is that a choice?

OpenStudy (anonymous):

your choices are missing some equal signs, so it is not clear but that is the right answer

OpenStudy (anonymous):

C is the right answer or the one you said. It doesn't have that as the answer. but it has y + 4 = 4 (x + 4)

OpenStudy (anonymous):

hmm hold on

OpenStudy (anonymous):

it definitely should be \[y+4=4(x-4)\]

OpenStudy (anonymous):

Alright it has that answer! I'm sorry!

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

then pick it!

OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through points S(–7,–6) and T(10,8). A. y = -14/17x + 4/17 B. y= -14/17x - -4/17 C. y = 14/17x - 4/17 D. y = 14/17x + 4/17 D is my answer.

OpenStudy (anonymous):

Write the equation for the horizontal line that contains point G(3, 4). Choose one answer. a. y = 4 b. y = 3 c. x = 3 d. x = 4 Answer: A Write the equation for the vertical line that contains point E(–7, 7). Choose one answer. a. y = –7 b. x = –7 c. y = 7 d. x = 7 Answer: B

OpenStudy (anonymous):

once again...

OpenStudy (anonymous):

A is right

OpenStudy (anonymous):

B is right too

OpenStudy (anonymous):

these are the last questions.

OpenStudy (anonymous):

that was it?